# Local Loss Optimization in the Infinite Width: Stable Parameterization of Predictive Coding Networks and Target Propagation

### v1

@@ 4.1 $\mu$P of PC with single-shot sequential inference
Figure 1: $\mu$P enables the transfer of learning rates across widths. (Left) PC reduces to SGD, when F-ini, FPA and SI are applied. In fact, using the $\mu$P of SGD, learning rates are successfully transferred across different widths. (Right) Even without FPA, our $\mu$P of PC also allows $\mu$Transfer across widths. In this case, inference is performed only once and the difference in test accuracy between $\bar{\gamma}_{L}=0$ and $\bar{\gamma}_{L}=1$ is small. Both figures show results for training a 3-layer MLP on FashionMNIST.
As noted in the preliminary section, PC involves such techniques as F-ini, SI, and FPA, which must be clearly distinguished when deriving the $\mu$P. It is well-established that when F-ini, SI, and FPA are all assumed, PC reduces to the gradient computation of BP, and the $\mu$P should match the one from standard BP. Figure 1 shows that when F-ini, FPA, and SI are applied, the $\mu$P of BP can be directly transferred to PC.
However, this may not hold for general PC and BP as there is no guarantee of their equivalence. To explore this, we first remove FPA. Although initialization (F-ini) and sampling (SI) are inherently arbitrary, the justification for FPA is unclear from both machine learning and biological perspectives. In PC without FPA, we find the $\mu$P as follows:
@@
Let the inference step sizes be $\gamma_{l<L}=\Theta(1)$ and $\gamma_{L}=\gamma^{\prime}/M^{\bar{\gamma}_{L}}$ with a positive constant $\gamma^{\prime}$. Consider the first one-step update of the learning parameters after a first single-shot SI with F-initialization. Then, PC admits the $\mu$P for feature learning at
@@
Rough sketch of the derivation. Section A.2 of the Appendix presents a detailed and comprehensive derivation. It is based on the perturbation approach, which applies to general networks with nonlinear activation functions. This method is inspired by the previous work that derived the $\mu$P by evaluating Conditions 3.1 and 3.2 using the perturbations, such as $\partial_{\eta^{\prime}}(\Delta W_{l,1}h_{l-1,1})\big{|}_{\eta^{\prime}=0}=\Theta(1)$. This allows for a systematic and transparent derivation. In PC, we extend the perturbation argument to the inference step size and require
@@
which is an example of Condition 3.1 for the hidden layer. Under the assumption of F-ini ($e_{l,0}=0$), by putting $\delta_{l}=\nabla_{u_{l}}\mathcal{L}$, we obtain $u_{l,1}-u_{l,0}=-\prod_{i=l+1}^{L}\gamma_{i}\delta_{l}$, and
@@
For the hidden layers, this leads to
$\partial_{\gamma^{\prime}}\partial_{\eta^{\prime}}(\Delta W_{l,1}h_{l-1,1})\big{|}_{\eta^{\prime}=\gamma^{\prime}=0}$ $=\frac{1}{M^{\theta_{l}+\bar{\gamma}_{L}}}\left(-\delta_{l}+\phi^{\prime}(W_{l}u_{l-1,0})\circ W_{l}\delta_{l-1}\right)h_{l-1,0}^{\top}h_{l-1,0}.$ (14)
We can similarly evaluate the other layers. Eventually, the $\mu$P’s conditions become
$\theta_{1}+\bar{\gamma}_{L}+(a_{L}+b_{L})=0,\quad\theta_{l}+\bar{\gamma}_{L}-1+(a_{L}+b_{L})=0,$ (15)
$\theta_{L}-1=0,\quad a_{L}+b_{L}-1=0.$ (16)
The last condition comes from Condition 3.2. We can derive the NTK parameterization of PC in the same way.
As Figure 1 demonstrates, the obtained $\mu$P supports $\mu$Transfer in PC without FPA. This means that the optimal hyperparameters tuned for smaller-width models can be effectively re-used in larger-width models. Additionally, consistent with previous work, we observed the empirical rule of “wider is better” in $\mu$P [Yang et al., 2021], where test accuracy improves as the network width increases.
@@ 4.2 Analysis with Linear Network
Figure 2: (Left) Comparison of gradients with the analytical solution of a linear network. (a) $\bar{\gamma}_{L}=0$ yields gradients closer to BP gradient (which means GD in this experiment) compared to $\bar{\gamma}_{L}=-1$. (b) As $M_{L}$ approaches 1, PC’s gradient converges to BP’s. (c) As $M_{l}$ increases, the PC gradient approaches BP’s. (Right) In a nonlinear MLP, PC’s gradient also approaches BP’s when $\bar{\gamma}_{L}=0$. This experiment was conducted using MLP with Tanh on the FashionMNIST dataset.
In the previous section, we derived the $\mu$P under the assumption that inference is performed only once using F-ini and SI. However, in practice, the inference phase typically involves multiple update steps. To address this, we found that it is possible to explicitly derive the following general solutions (fixed points) of the inference phase for linear networks. See Section A.3 for the derivation.
@@
Suppose an L-layered linear network, where we put $e_{l}^{*}=v_{l}^{*}-W_{l}v_{l-1}^{*}$, with * denoting the fixed point of the inference process(3). The following holds:
@@
$v_{l}^{*}$ $=W_{l:1}x+\left(\frac{\gamma_{L}}{\gamma_{l}}W_{L:l+1}^{\top}+\sum_{i=2}^{l-1}\frac{\gamma_{L}}{\gamma_{i-1}}W_{l:i}W_{L:i}^{\top}\right)(I+C_{\gamma}(W))^{-1}(y-f).$ (18)
@@
Suppose the setting of Theorem 4.2, $\gamma_{l<L}=\Theta(1)$ and $a_{L}+b_{L}=1$. In the infinite-width limit, the PC’s gradient reduces to GD for $\gamma_{L}=\Theta(1)$. For $\gamma_{L}=\Theta(M)$, the preconditioner part $C_{\gamma}$ remains of order 1.
@@
The exact solutions $e_{l}^{*}$ provide much clearer insight into the gradient computation compared to Eq. (7), which was previously argued but not explicitly solved. First, it becomes evident that PC does not generally coincide with GNT. Consequently, PC is also generally different from TP. In fact, PC coincides with GNT only for the input layer in a shallow network (i.e., $L=2$), where the update vector for PC corresponds to a GNT update with a damping term. Although PC does not entirely coincide with GNT, it is noteworthy that the scaling of $c_{l}$ in $\mu$P for $\bar{\gamma_{L}}=-1$ matches that of GNT. In contrast, for $\bar{\gamma}_{L}=0$, the PC’s gradient aligns with the (S)GD. Because the preconditioner part scales as $C_{\gamma}=O(1/M)$ in the infinite-width limit, we observe that $e_{l}^{*}=\delta_{l}$, which reduces to the GD. Naturally, $\mu$P matches that of GD in this case.
@@
Second, the order of $e_{l}^{*}$ in the analytical solution for the linear network matches the order of $e_{l,1}$ as derived in Theorem 11. Therefore, this theorem implies that in linear networks, the $\mu$P of PC would remains unchanged regardless of the presence of F-ini or the number of inference iterations. Moreover, as proved in Section A.3.2, the orders of $e_{l}^{*}$ and $e_{l,1}$ align only when $\gamma_{l<L}=\Theta(1)$. In practical settings with multiple inferences, it is desirable for the $\mu$P to be consistent both when performing a single inference and after the inference has fully converged. Therefore, setting $\gamma_{l<L}=\Theta(1)$ is reasonable.
@@
Additionally, we found that the dimension of the last layer plays a key role in determining the similarity between PC and BP. According to the solution for linear networks, when $M_{L}=1$, the PC’s gradient aligns with GD. Figure 2 shows numerical results confirming that for $M_{L}=1$, the gradient direction always corresponds to GD, and for $M_{l}\gg M_{L}=\Theta(1)$, the gradient approaches GD as well. We observed that both GN and BP get much closer to each other for sufficiently large widths. In other words, even when we realize GNT by setting $\bar{\gamma}_{L}=-1$, it has a quite close direction. A detailed view of the cosine similarity at the large width is shown in Figure 2(c). This result seems reasonable because in the context of second-order optimization, it has also been reported that GNT tends to collapse into an identity matrix owing to damping [Benzing, 2022]. In summary, while PC’s gradient switches between GD and GNT depending on the parameterization, it is important to highlight that GNT behaves similar to GD in the infinite-width limit.
@@
As a minor extension, we can also analyze the nudge-type loss of PC defined by Eq. (S.86)  [Alonso et al., 2022, Millidge et al., 2023, Pinchetti et al., 2024]. In this case, the damping term $I$ in Eq. (17), is replaced by $(1+\gamma/\beta)I$. Thus, the dependence on the parameterization remains essentially the same as that of the naive PC. Further discussion on nudge-type PC can be found in Appendix A.3.3.
@@ 4.3 Stability of inference phase
Figure 3: (Left) $\hat{\gamma_{L}}=-1$ steadily reduces the local loss as width increases. We observed the inference loss in a randomly initialized linear network for various $\bar{\gamma}_{L}$. For $\bar{\gamma}_{L}=-1$, the inference loss consistently decreases with increasing width. (Right) The "wider is better" trend holds for $\mu$P with $\hat{\gamma_{L}}=-1$. With F-ini, this trend holds for $\mu$P regardless of the $\bar{\gamma}_{L}$ value. However, without F-ini, the benefits of $\bar{\gamma}_{L}=-1$ become particularly prominent.
@@
Figure 4: (Left) $\mu$P can transfer the learning rate across widths (without F-ini). We trained a 3-layer CNN on FashionMNIST with 100 inference iterations. Without F-ini, the stability of the inference becomes more crucial. As a result, unlike the single-shot SI with F-ini shown in Figure 1, the stability provided by $\bar{\gamma}_{L}=-1$ becomes critical. Note that additional experiments with different settings, including other architectures and loss functions, are presented in Section B.1.2 of the Appendix. (Right) $\Delta h$ remains consistent across widths during training. We confirm that the condition $\Delta h=\Theta(1)$ required by $\mu$P holds throughout the training.
@@
To ensure feature learning in SGD, the $\mu$P framework requires stable activations, i.e., $\Delta u_{l<L}=\Theta(1)$. It seems natural to apply this requirement to the inference phase of PC. That is, let us suppose $u_{l<L,s}$ varies by $\Theta(1)$ during the inference. Note that in Eq.(3) at $L-1$, the feedforward signal from the lower layer is $\gamma_{L-1}e_{L-1,s}=\Theta(1)$, and the error feedback from the last layer is $\gamma_{L}\phi^{\prime}\left(u_{L-1,s}\right)\circ W_{L}^{\top}e_{L,s}=\Theta\left(1/M^{\bar{\gamma}_{L}+b_{L}-1}\right)$. Both terms should be of order $\Theta(1)$ for the inference to successfully merge both feedforward and feedback signals. When $b_{L}=1$, this condition requires $\gamma_{L}=\Theta(M)$, and we can expect the local loss in the last layer $e_{L}$ to decrease most prominently during the inference. Additionally for $\gamma_{l<L}=\Theta(1)$, the inference remains stable for layers $l<L-1$. Empirical results in Figure 3 (left) comfirm that when $\bar{\gamma}_{L}=-1$, the inference loss decreases consistently as the width increase, verifying that the “wider is better” hypothesis holds even in inference. This facilitates the hyperparameter transfer of $\gamma_{L}$ for the inference dynamics.
We also observe the benefits of using $\bar{\gamma}_{L}=-1$ for the parameter updates. Without F-ini, the convergence of inference usually deteriorates for SP, making inference stability specially critical in this scenario. As shown in Figure 3 (right), the “wider is better” trend holds with F-ini regardless of $\bar{\gamma}_{L}$. However, without F-ini, this trend holds only when $\bar{\gamma}_{L}=-1$. Figure 4 demonstrates that the $\mu$Transfer of the learning rate holds for $\bar{\gamma}_{L}=-1$. Additionally, Figure 5 indicates that $\bar{\gamma}_{L}=-1$ is also preferable from the perspective of $\mu$Transfer of $\gamma_{L}$.

### v2

@@ 4.1 $\mu$P of PC with single-shot sequential inference
Figure 1: $\mu$P enables the transfer of learning rates across widths. (Left) PC reduces to SGD when F-ini, FPA, and SI are applied. In fact, using the $\mu$P of SGD, learning rates are successfully transferred across different widths. (Right) Even without FPA, our $\mu$P of PC also allows $\mu$Transfer across widths. In this case, inference is performed only once, and the difference in test accuracy between $\bar{\gamma}_{L}=0$ and $\bar{\gamma}_{L}=1$ is small. Both figures show results with a 3-layer MLP on FashionMNIST.
As noted in section 3.1.1, PC involves such techniques as F-ini, FPA and SI, which must be clearly distinguished when deriving the $\mu$P. It is well-established that when F-ini, SI, and FPA are all assumed, PC reduces to the gradient computation of BP, and the $\mu$P matches that of standard BP. Figure 1 shows that when F-ini, FPA, and SI are applied, the $\mu$P of BP can be directly transferred to PC and leads to learning rate transfer across width. However, this may not hold for general PC and BP as there is no guarantee of their equivalence. To explore this, we first remove FPA. Although initialization (F-ini) and sampling (SI) are inherently arbitrary, the justification for FPA is unclear from both machine learning and biological perspectives. In PC without FPA, we find the $\mu$P as follows:
@@
Let the inference step sizes be $\gamma_{l<L}=\Theta(1)$ and $\gamma_{L}=\gamma^{\prime}/M^{\bar{\gamma}_{L}}$ with a positive constant $\gamma^{\prime}$. Consider the first one-step update of the learning parameters after a first single-shot SI with F-ini. Then, PC admits the $\mu$P for feature learning at
@@
Rough sketch of the derivation. Section B.1 of the Appendix presents a detailed and comprehensive derivation. It is based on the perturbation approach, which applies to general networks with nonlinear activation functions. This method is inspired by the previous work that derived the $\mu$P by evaluating Conditions 3.1 and 3.2 using the perturbations, such as $\partial_{\eta^{\prime}}(\Delta W_{l,1}h_{l-1,1})\big{|}_{\eta^{\prime}=0}=\Theta(1)$. This allows for a systematic and transparent derivation. In PC, we extend the perturbation argument to the inference step size and require
@@
which is an example of Condition 3.1 for the hidden layer. Under the assumption of F-ini ($e_{l,0}=0$), by putting $\delta_{l}=\nabla_{u_{l}}u_{L}$ ($l<L$) and $\delta_{L}=y-W_{L}v_{L-1,0}$, we obtain $u_{l,1}-u_{l,0}=-\prod_{i=l+1}^{L}\gamma_{i}\delta_{l}$, and
@@
For the hidden layers, the perturbation term (13) becomes $M^{-(\theta_{l}+\bar{\gamma}_{L})}(-\delta_{l}+\phi^{\prime}(W_{l}u_{l-1,0})\circ W_{l}\delta_{l-1})h_{l-1,0}^{\top}h_{l-1,0}$ and we obtain $\theta_{l}+\bar{\gamma}_{L}-1+(a_{L}+b_{L})=0$. We can similarly evaluate the other layers. The last condition, $b_{L}=1$, comes from Condition 3.2. We can derive the NTK parameterization of PC in the same way.
As Figure 1 demonstrates, the obtained $\mu$P supports $\mu$Transfer in PC without FPA. Note that $\mu$Transfer is defined as satisfying both conditions: the optimal learning rate can be set independently of the order of width, and the empirical rule that ‘wider is better’ holds (Yang et al., 2021). This means that the optimal hyperparameters tuned for smaller-width models can be effectively re-used in larger-width models. Additionally, consistent with previous work, we observed the empirical rule of “wider is better” in $\mu$P (Yang et al., 2021), where test accuracy improves as the network width increases. The derivation of $\mu$P through a one-step update can be immediately generalized to cross-entropy loss in the same way as for $\mu$P of naive gradient descent. Thus, $\mu$Transfer can similarly be observed for cross-entropy loss, as shown in Appendix (Figure S.5).
@@ 4.2 Analysis with Linear Network
Figure 2: (Left) Comparison of gradients with the analytical solution of a linear network. We measured the cosine similarity between the gradients analytically derived in Theorem 4.2 and the BP gradients or GN gradients for each layer. (a) As $M_{L}$ approaches 1, PC’s gradient converges to BP’s. (b) As $M_{l}$ increases, the PC gradient approaches BP’s. (c) $\bar{\gamma}_{L}=0$ yields gradients closer to BP gradient (which means SGD in this experiment) compared to $\bar{\gamma}_{L}=-1$. (Right) In a nonlinear MLP, PC’s gradient also approaches BP’s when $\bar{\gamma}_{L}=0$.
In the previous section, we derived the $\mu$P under the assumption that inference is performed only once using F-ini and SI. However, in practice, the inference phase typically involves multiple update steps. To address this, we found that it is possible to explicitly derive the following general solutions (fixed points) of the inference phase for linear networks. See Section B.2 for the derivation.
@@
Suppose an L-layered linear network and a mean squared error loss $\mathcal{L}(y,W_{L}v_{L-1})$, and put $e_{l}^{*}=v_{l}^{*}-W_{l}v_{l-1}^{*}$, with * denoting the fixed point of the inference process (3). The following holds:
@@
$v_{l}^{*}$ $=W_{l:1}x+(\frac{\gamma_{L}}{\gamma_{l}}W_{L:l+1}^{\top}+\sum_{i=2}^{l-1}\frac{\gamma_{L}}{\gamma_{i-1}}W_{l:i}W_{L:i}^{\top})(I+C_{\gamma}(W))^{-1}(y-f).$ (16)
@@
Suppose the setting of Theorem 4.2, $\gamma_{l<L}=\Theta(1)$ and the random weights given by $\mu$P i.e., $a_{L}+b_{L}=1$. In the infinite-width limit, the PC’s gradient reduces to the first-order GD for $\gamma_{L}=\Theta(1)$. For $\gamma_{L}=\Theta(M)$, the preconditioner part $C_{\gamma}$ remains of order 1.
@@
The exact solutions $e_{l}^{*}$ provide much clearer insight into the gradient computation compared to Eq. (8), which was previously argued but not explicitly solved. First, it becomes evident that PC does not generally coincide with GNT. Consequently, PC is also generally different from TP. In fact, PC coincides with GNT only for the input layer in a shallow network (i.e., $L=2$), where the update vector for PC corresponds to a GNT update with a damping term. Although PC does not entirely coincide with GNT, it is noteworthy that the scaling of $c_{l}$ in $\mu$P for $\bar{\gamma_{L}}=-1$ matches that of GNT. In contrast, for $\bar{\gamma}_{L}=0$, the PC’s gradient aligns with the first-order GD. Because the preconditioner part scales as $C_{\gamma}=O(1/M)$ in the infinite-width limit, we observe that $e_{l}^{*}=\delta_{l}$, which reduces to the first-order GD. Naturally, $\mu$P matches that of first-order GD in this case. Intuitively, $\gamma_{L}$ reflects how effectively the last layer’s error propagates downward. Like linear regression, the last layer’s inference solution inherently involves an inverse matrix. Thus, when $\gamma_{L}$ is larger than other $\gamma_{l}$ values, the last layer’s representation is computed first and propagated downward, making the solution resemble GNT.
@@
Second, the order of $e_{l}^{*}$ in the analytical solution for the linear network matches the order of $e_{l,1}$ as derived in Theorem 12. Therefore, this theorem implies that in linear networks, the $\mu$P of PC would remain unchanged regardless of the presence of F-ini or the number of inference iterations. Moreover, as proved in Section B.2.3, the orders of $e_{l}^{*}$ and $e_{l,1}$ align only when $\gamma_{l<L}=\Theta(1)$. In practical settings with multiple inferences, it is desirable for the $\mu$P to be consistent both after a single inference and after the inference has fully converged. Therefore, setting $\gamma_{l<L}=\Theta(1)$ is reasonable.
@@
Additionally, we found that the dimension of the last layer plays a key role in determining the similarity between PC and BP. According to the solution for linear networks, when $M_{L}=1$, the PC’s gradient aligns with GD. Figure 2 shows numerical results confirming that for $M_{L}=1$, the gradient direction always corresponds to GD, and for $M_{l}\gg M_{L}=\Theta(1)$, the gradient approaches GD as well. We observed that both GN and BP get much closer to each other for sufficiently large widths. In other words, even when we realize GNT by setting $\bar{\gamma}_{L}=-1$, it has a quite close direction. A detailed view of the cosine similarity at the large width is shown in Figure 2(c). This result seems reasonable because in the context of second-order optimization, it has also been reported that GNT tends to collapse into an identity matrix owing to damping (Benzing, 2022). In summary, while PC’s gradient switches between first-order GD and GNT depending on the parameterization, it is important to highlight that GNT behaves similarly to GD in the infinite-width limit.
@@
As a minor extension, we can also analyze the nudge-type loss of PC defined by Eq. (S.10)  (Alonso et al., 2022; Millidge et al., 2023; Pinchetti et al., 2024). In this case, the damping term $I$ in Eq. (15), is replaced by $(1+\gamma/\beta)I$. Thus, the dependence on the parameterization remains essentially the same as that of the naive PC. Further discussion on nudge-type PC can be found in Appendix B.2.4.
@@ 4.3 Stability of inference phase
Figure 3: (Left) $\bar{\gamma_{L}}=-1$ steadily reduces the local loss as width increases. We observed the inference loss in a randomly initialized linear network for various $\bar{\gamma}_{L}$. For $\bar{\gamma}_{L}=-1$, the inference loss consistently decreases with increasing width. (Right) The "wider is better" trend holds for $\mu$P with $\bar{\gamma_{L}}=-1$. With F-ini, this trend holds for $\mu$P regardless of the $\bar{\gamma}_{L}$ value. However, without F-ini, the benefits of $\bar{\gamma}_{L}=-1$ become particularly prominent.
@@
Figure 4: (Left) $\mu$P can transfer the learning rate across widths (without F-ini). We trained a 3-layer CNN on FashionMNIST with 100 inference iterations. Without F-ini, the stability of the inference becomes more crucial. As a result, unlike the single-shot SI with F-ini shown in Figure 1, the stability provided by $\bar{\gamma}_{L}=-1$ becomes critical. Note that additional experiments under different settings, including those with VGG5 (Figure S.3) and cross-entropy loss (Figure S.5), are presented in Section D.1.2 of the Appendix. (Right) $\Delta h$ remains consistent across widths during training. We confirm that the condition $\Delta h=\Theta(1)$ required by $\mu$P holds throughout the training.
@@
To ensure feature learning in SGD, the $\mu$P framework requires stable activations, i.e., $\Delta u_{l<L}=\Theta(1)$. It seems natural to apply this requirement to the inference phase of PC. That is, let us suppose $u_{l<L,s}$ varies by $\Theta(1)$ during the inference. Note that in Eq.(3) at $L-1$, the feedforward signal from the lower layer is $\gamma_{L-1}e_{L-1,s}=\Theta(1)$, and the error feedback from the last layer is $\gamma_{L}\phi^{\prime}\left(u_{L-1,s}\right)\circ W_{L}^{\top}e_{L,s}=\Theta\left(1/M^{\bar{\gamma}_{L}+b_{L}}\right)$. Both terms should be of order $\Theta(1)$ for the inference to successfully merge both feedforward and feedback signals. When $b_{L}=1$, this condition requires $\gamma_{L}=\Theta(M)$, and we can expect the local loss in the last layer $e_{L}$ to decrease most prominently during the inference. Additionally for $\gamma_{l<L}=\Theta(1)$, the inference remains stable for layers $l<L-1$. Empirical results in Figure 3 (left) confirm that when $\bar{\gamma}_{L}=-1$, the inference loss decreases consistently as the width increases, verifying that the “wider is better” hypothesis holds even in inference. This facilitates the hyperparameter transfer of $\gamma_{L}$ for the inference dynamics.
We also observe the benefits of using $\bar{\gamma}_{L}=-1$ for the parameter updates. Without F-ini, the convergence of inference usually deteriorates for SP, making inference stability especially critical in this scenario. As shown in Figure 3 (right), the “wider is better” trend holds with F-ini regardless of $\bar{\gamma}_{L}$. However, without F-ini, this trend holds only when $\bar{\gamma}_{L}=-1$. Figure 4 demonstrates that the $\mu$Transfer of the learning rate holds for $\bar{\gamma}_{L}=-1$. Additionally, Figure 5 indicates that $\bar{\gamma}_{L}=-1$ is also preferable from the perspective of $\mu$Transfer of $\gamma_{L}$.